Hello, I've tried to solve this problem t^2/3=4 and my solution was 8 because I put (t^2/3) ^3/2 = 4^3/2 and t= (square root)4^3 = (square root) 64 = 8. Can you help me please? Thank You

Answer by checkley77(12844)

t^2/3=4
cubert(t^2)=4
t^2=4^3
t^2=64
t=sqrt64
t=8 ans.
Proof:
8^2/3=4
cubert8^2=4
cubert64=4
4=4
---------------------
There are usually more than 1 way to solve math problems.
Pick the 1 that you can remember & hopefully it will be the simplest & minimal chance to make errors.
ALWAYS do the proof.